solve equation 3|3x-4|-7=5
Answer by robertb(5830)

3|3x - 4| = 12
==> |3x - 4| = 12/3 = 4
==> |3x - 4| = 4
==> 3x - 4 = 4 or 3x-4 = -4
==> 3x = 8 or 3x = 0
==> x = 8/3 or x = 0.